OpenStudy (littleheron):

Will medal and fan! Solve for y. Show all your work. 6x - 2y = 24

OpenStudy (anonymous):

Let's solve for x. 6x−2y=24 Step 1: Add 2y to both sides. 6x−2y+2y=24+2y 6x=2y+24 Step 2: Divide both sides by 6. 6x 6 = 2y+24 6 x= 1 3 y+4

OpenStudy (anonymous):

Now join my fan army :D

OpenStudy (littleheron):

it says i have been blocked

OpenStudy (anonymous):

? lol when and how?

OpenStudy (anonymous):

I unblocked you :D sorry about that